OpenStudy (anonymous):

√(5a+2)=-√(a+7) I got a= -3/2 please check my answer

myininaya (myininaya):

\[\sqrt{5a+2}=-\sqrt{a+7}\] square both sides \[5a+2=a+7\] \[5a-a=7-2\] \[4a=5\] \[a=\frac{4}{5}\]

myininaya (myininaya):

but there is no solution

OpenStudy (anonymous):

If 4a=5 then a= 5/4

myininaya (myininaya):

yep i made a mistake there but still there is no solution